OpenStudy (anonymous):

Carpet is sold in square yards. How much carpet would a person need to buy for a rectangular room that has dimensions of 10 yards by 18 feet? 180 square yards 60 square yards 540 square yards 15 square yards

OpenStudy (anonymous):

This one also, you have to make the units equal to each other. 1ft = 1/3 yd 18ft = 6 yd

OpenStudy (anonymous):

First, we need to knwo the conversion between yards and feet. How many feet are in a yard?

OpenStudy (anonymous):

Then, you just simply multiply the sides together 10*6 = 60 yd^2

OpenStudy (anonymous):